Please Help Me!!!
The Length of a rectangle is 5in more than the width. if the perimeter is 70in, what is the width?
Please could someone kindly help me
Thanks

Answer by rfer(16322)

2x+2(x+5)=70
2x+2x+10=70
4x=60
x=15
x+5=20